[Brahms-l] Shift report 20050512 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu May 12 2005 - 00:11:37 EDT
Shifter leader was Thomas Materna
Shifters were 


Machine setup until 30 min before end of shift. Loss of yellow beam intensity, beam dumped twice.
runs 14584 to 14591.
No problems with detector, just a little problem with D2 magnet at the beginning of the shift.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 12 00:11:54 2005

This archive was generated by hypermail 2.1.8 : Thu May 12 2005 - 00:12:06 EDT